OpenStudy (anonymous):

Shortest distance between the two lines and find two points (one on each line) which are closest together. [1,0,−1]^T +t[1,1,1]^T and [0,1,1]^T +s[2,2,2]^T

OpenStudy (anonymous):

@phi help!

OpenStudy (phi):

these lines are parallel (same direction vector (1,1,1) )

OpenStudy (anonymous):

can you please show me how you got the answer

OpenStudy (phi):

I did not get the answer. I just pointed out that (1,1,1) and (2,2,2) point in the same direction, so the lines are parallel. we still have to figure out how far apart they are.

OpenStudy (anonymous):

ohh i see

OpenStudy (anonymous):

so there is no two points closest to each other

OpenStudy (phi):

there are lots of points in nearest to each other. I think you need to find one pair, and the distance.

OpenStudy (anonymous):

how do i find the distance

OpenStudy (phi):

ok, I finally had a thought.

OpenStudy (phi):

 both lines are going in the same direction. we can use that direction vector as a normal to a plane in 3d. For example, if we look at the plane that contains point [0,1,1] and has a normal vector (1,1,1), we have this equation \[ P \cdot (1,1,1) = b\] replacing P with (0,1,1) we find b is 2, and the equation of the plane is \[ P \cdot (1,1,1) = 2\] now we need the point on the other line that is in this same plane Let P= \( (1,0,−1)^T +t(1,1,1)^T \) \[ ( (1,0,−1)^T +t(1,1,1)^T) \cdot (1,1,1) = 2 \\ (1,0,−1)\cdot (1,1,1)+ t (1,1,1)\cdot (1,1,1)= 2 \\ 0 + 3t=2\\ t= \frac{2}{3} \] that means the point on the other line is (1,0,−1) +(⅔)(1,1,1) = (5/3 , ⅔, -⅓)